# Komaba, Meguro-ku, Tokyo 153-8902, Japan

### v2

@@ Abstract
We study perturbative unitarity bounds on the field-space curvature in de Sitter spacetime, using the momentum-space entanglement approach recently proposed by Pueyo, Goodhew, McCulloch, and Pajer. As an illustration, we perform a perturbative computation of the purity in two-scalar models and compare the resulting unitarity bounds with those obtained via a flat space approximation. In particular, we find that perturbative unitarity imposes an upper bound on the field-space curvature of the Hubble scale order, in addition to a bound analogous to the flat space result. This reflects the thermal nature of de Sitter spacetime. We also discuss generalizations to higher-dimensional field spaces.
@@ 1 Introduction
Notably, EFT becomes more powerful when combined with the concept of ultraviolet (UV) completion and fundamental principles such as unitarity. In particular, the S-matrix unitarity offers a powerful criterion for quantifying the UV cutoff scale and searching for new physics required for UV completion. Historically, the Higgs boson was predicted to restore unitarity in the high-energy scattering of weak bosons lee1991weak ; Lee:1977yc ; Dicus:1973gbw ; Chanowitz:1985hj . Similarly, string theory, as a UV-complete theory of gravity, emerged from S-matrix theory Veneziano:1968yb . Moreover, the study of UV completion has in turn revealed that not every EFT is UV completable, leading to nontrivial UV constraints on low-energy effective theories (see, e.g., deRham:2022hpx for a review article).
@@
While the S-matrix unitarity offers a fundamental tool to theoretically connect various scales in nature, cosmology is beyond its scope because scattering amplitudes are not well-defined in cosmological backgrounds. To address this limitation, several approaches have been taken so far: A pragmatic approach is to employ the flat space approximation and directly apply implications of the S-matrix unitarity to cosmological models Baumann:2015nta ; Kim:2019wjo ; Grall_2020 ; Melville_2020 ; Kim:2021pbr ; Freytsis:2022aho ; Grall:2021xxm . While this approach has limitations in applicability, the approximation can be justified at least physically as long as our focus is, e.g., on effective couplings generated by the UV dynamics well above the Hubble scale or, in other words, by the dynamics well inside the horizon.
@@
A more challenging direction would be to set up the bootstrap program in cosmology, motivated by recent progress in the S-matrix bootstrap and the conformal bootstrap. Aiming at this ultimate goal, unitarity and (non-)analyticity of cosmological correlators and wavefunctions of the universe have been studied intensively under the slogan of the Cosmological Bootstrap Arkani-Hamed:2018kmz ; Baumann:2019oyu ; Baumann:2020dch ; Arkani-Hamed:2017fdk ; Benincasa:2018ssx ; Sleight:2019mgd ; Sleight:2019hfp ; Goodhew:2020hob ; Cespedes:2020xqq ; Pajer:2020wxk ; Jazayeri:2021fvk ; Bonifacio:2021azc ; Melville:2021lst ; Goodhew:2021oqg ; Pimentel:2022fsc ; Jazayeri:2022kjy ; Qin:2022fbv ; Xianyu:2022jwk ; Wang:2022eop ; Qin:2023ejc ; Stefanyszyn:2023qov ; DuasoPueyo:2023kyh ; Cespedes:2023aal ; Bzowski:2023nef ; Arkani-Hamed:2023kig ; Grimm:2024mbw ; Aoki:2024uyi ; Stefanyszyn:2024msm ; Liu:2024xyi ; Goodhew:2024eup ; Ghosh:2024aqd ; Lee:2024sks ; Cespedes:2025dnq ; Pimentel:2025rds ; Stefanyszyn:2025yhq ; Qin:2025xct . There are also alternative attempts to define the notion of the S-matrix itself to cosmological backgrounds Mack:2009mi ; Penedones:2010ue ; Marolf:2012kh ; Melville:2023kgd ; Donath:2024utn ; Melville:2024ove ; Taylor:2024vdc ; Ferrero:2021lhd ; Mandal:2019bdu ; Spradlin:2001nb ; Mei:2024sqz . These developments motivate further studies toward refinement of the S-matrix unitarity as a guiding principle in cosmology.
@@
Building upon this insight, an interesting approach was proposed recently in Pueyo:2024twm to utilize entanglement measures such as purity and entanglement entropy to derive unitarity constraints on cosmological models (see also Colas:2022kfu ; Colas:2024xjy ; Burgess:2024eng ; Ueda:2024cyf ; Balasubramanian:2011wt ; Aoude:2024xpx ; Cheung:2023hkq ; Peschanski:2016hgk ; Kowalska:2024kbs ; Brahma:2023lqm ; Boutivas:2023mfg for related developments). A key of this approach is in the fact that interactions induce entanglement between momentum modes, leading to an analogy between entanglement measures and scattering amplitudes. Crucially, this framework is applicable even in curved spacetime as long as the density matrix is well-defined. Based on this approach, perturbative unitarity bounds in inflationary backgrounds were studied in particular.
@@
In this paper, we apply the momentum-space entanglement approach of Pueyo:2024twm to derive perturbative unitarity bounds on field-space curvature of nonlinear sigma models, which widely appear for example as effective theories of (pseudo-)Nambu-Goldstone bosons, in de Sitter spacetime. Unlike the original paper, we study the unitarity bounds on purity without taking the superhorizon limit, which allows us to perform detailed analysis of the bounds that interpolate the flat space analysis and the superhoziron analysis. Interestingly, we find that the perturbative unitarity gives an upper bound on the field space curvature of the Hubble scale order, in addition to a bound similar to the flat space result, reflecting the thermal nature of de Sitter spacetime.
@@
This paper is organized as follows: In Sec. 2, we review the momentum-space entanglement approach to perturbative unitarity proposed in Pueyo:2024twm . In particular, we introduce a perturbative formula of purity and its uitarity condition. In Sec. 3, we study the perturbative unitarity bounds in flat spacetime and show that the UV cutoff is set by the field-space curvature, similarly to the bounds obtained from scattering amplitudes. In Sec. 4, we extend the analysis to de Sitter spacetime. In addition to a bound similar to the flat space result, we find an upper bound on the field-space curvature at the order of the Hubble scale. We conclude our analysis in Sec. 5. Technical details are collected in the Appendices.
@@
Throughout the paper, we adopt the metric signature (,+,+,+)(-,+,+,+)( - , + , + , + ), Greek letters $\mu,\nu,\cdots$ to denote spacetime indices, and shorthand notations for the integration measure,
@@ 2 Perturbative unitarity bounds from purity: a brief review
This section gives a brief review of the momentum-space entanglement approach to perturbative unitarity bounds proposed in Pueyo:2024twm . In particular, we consider purity that quantifies the entanglement between a system of interest and its complement (environment). In Sec. 2.1, we first introduce the concept of purity in quantum field theory (QFT) and briefly explain how it can be evaluated using the wavefunction representation. Then, in Sec. 2.2, we consider multi-scalar models with nonzero field-space curvature and provide a concrete formula for the purity, which is used in the following sections.
@@ 2.1 Momentum-space entanglement and purity in EFT
Consider a quantum system in a pure state represented by a density matrix111 In Pueyo:2024twm , the density matrix was defined without being canonically normalized to carefully discuss regularization for the continuum and infinite-volume limit. The density matrix here should also be understood under the same regularization, even though we assume canonical normalization for visual clarity.,
$\rho=\ket{\Omega}\bra{\Omega}\,,\quad{\operatorname{Tr}}\,\rho=1\,.$ (2)
@@
Next, we consider momentum-space entanglement in QFT. For this, it is convenient to employ the field eigenstate |$\ket{\phi}$ as a basis of the Hilbert space and express the density matrix $\rho$ at a given time $\eta=\eta_{*}$ as
$\rho$ =|$=\ket{\Omega}\bra{\Omega}$ |
@@
$=\int\mathcal{D}\phi\mathcal{D}\phi^{\prime}\rho_{\phi\phi^{\prime}}\ket{\phi}\bra{\phi^{\prime}}\,,$ (5)
where we define the components of the density matrix as $\rho_{\phi\phi^{\prime}}\coloneqq\braket{\phi}{\Omega}\braket{\Omega}{\phi^{\prime}}$. Here and in the rest of this subsection, we focus on a single real scalar model for illustration, but its extension to general setups is straightforward.
@@
Its violation signals breakdown of the EFT, hence we can use the purity bound (9) to identify the maximum energy range of validity of the EFT222 Alternatively, if the energy scale of interest is specified and the cutoff scales $\Lambda_{\text{IR}}$ and $\Lambda_{\text{UV}}$ are given, one can interpret the bounds (9) as consistency conditions on the model parameters such as the particle spectrum and the coupling constants. .
@@
In practical computations of the purity, it is convenient to introduce the Schrödinger wave functional $\Psi[\phi]$, which is defined by the inner product of the state |$\ket{\Omega}$ and the field eigenstate |$\ket{\phi}$ at the time $\eta=\eta_{*}$ as
$\Psi[\phi]\coloneqq\braket{\phi}{\Omega}\,.$ (10)
@@
where $\phi_{{\bm{k}}}$ is the Fourier mode of $\phi$ and the kernels $\psi_{n}({\bm{k}}_{1},\cdots{\bm{k}}_{n})$ are called wavefunction coefficients. We further assume translation invariance along the spatial directions of the theory and the state |$\ket{\Omega}$. Besides, an overall constant factor has been suppressed, as it does not affect the subsequent analysis.
@@ 2.2 Perturbative formula for purity
$S$ $=\int{\mathrm{d}}^{4}x\sqrt{-g}\left[-\frac{1}{2}\sum_{I}\Big{(}(\partial_{\mu}\phi^{I})^{2}+m_{I}^{2}(\phi^{I})^{2}\Big{)}+\sum_{I,J,K,L}\frac{1}{2}C_{IJKL}\phi^{K}\phi^{L}\partial_{\mu}\phi^{I}\partial^{\mu}\phi^{J}+\cdots\right]\,,$ (15)
where the dots stand for higher order terms in $\phi^{I}$ and they are irrelevant in the following analysis. We study this model in homogeneous and isotopic spacetime,
@@
More specifically, we consider flat spacetime in Sec. 3 and de Sitter spacetime in Sec. 4. As for the state |$\ket{\Omega}$, we consider the free theory vacuum for flat spacetime and the Bunch-Davies vacuum for de Sitter spacetime, respectively.
@@
×$\hskip 60.0pt\times\left(\partial_{\eta}K_{\phi^{I}}(k_{1};\eta)\partial_{\eta}K_{\phi^{J}}(k_{2};\eta)+({\bm{k}}_{1}\cdot{\bm{k}}_{2})K_{\phi^{I}}(k_{1};\eta)K_{\phi^{J}}(k_{2};\eta)\right)\Big{]}$
[…]
@@
Finally, we provide a formula for the purity. Let us choose the Fourier modes $\phi^{\bar{I}}_{{\bm{p}}}$ and $\phi^{\bar{I}}_{-{\bm{p}}}$ of the species label $\bar{I}$ as the system of interest. Then, from the definition (6)-(7) and the wavefunction representation of the density matrix (11), the purity can be evaluated at the tree level as333 See the original paper Pueyo:2024twm for details of the diagrammatic method for the purity computation.
@@
with $I_{\phi^{\bar{I}}}(\Lambda_{\text{IR}},\Lambda_{\text{UV}},{\bm{p}})$ given by444 $\phi^{\bar{I}}_{\pm{\bm{p}}}$ are in the system sector, so that the momentum integral over the environment modes of $\phi^{I}$ has to be performed such that ${\bm{k}}_{i}\neq{\bm{p}}$ to be precise. However, this gives a measure-zero effect and negligible, so that we do not care in the following analysis.
@@ 3 Perturbative unitarity bounds in flat spacetime
We begin with the flat spacetime, for which perturbative unitarity bounds on the field space curvature were already studied using scattering amplitudes (see, e.g., Nagai:2019tgi ). To be concrete, this section focuses on the following two-scalar model as a simple case of (15):
@@ 3.2.1 Bounds on $\gamma_{\phi}$
First, we consider $\gamma_{\phi}$, choosing the Fourier modes $\phi_{\bm{p}}$ and $\phi_{-{\bm{p}}}$ as the system. From the formula (21), the corresponding purity reads555 Thanks to the rotational symmetry, the momentum-dependence of the purity in our setup is only through the amplitude $p=|{\bm{p}}|$ of the momentum ${\bm{p}}$ of the system modes. Also, it is convenient for visual clarity to suppress its cutoff-dependence. Hence, we employ the notation $\gamma_{\phi^{\bar{I}}}(p)$ and $I_{\phi^{\bar{I}}}(p)$ for $\gamma_{\phi^{\bar{I}}}(\Lambda_{\text{IR}},\Lambda_{\text{UV}},{\bm{p}})$ and $I_{\phi^{\bar{I}}}(\Lambda_{\text{IR}},\Lambda_{\text{UV}},{\bm{p}})$ in the rest of the paper.
@@ 4.1 Illustrative example: $m_{\phi}=0$ and $m_{\sigma}=\sqrt{2}H$
×$\quad\times\bigg{[}k_{1}k_{2}\Big{(}1+(1-ik_{T}\eta_{*})^{2}\Big{)}(1-\cos\theta)-\Big{(}(k_{1}+k_{2})\left(1-ik_{T}\eta_{*}\right)+k_{T}\Big{)}k_{T}\cos\theta\bigg{]}\,,$ (50)
@@
Below, we present the bounds on $\gamma_{\phi}$ and $\gamma_{\sigma}$, and their implications in order.

### v3

@@ Abstract
We study perturbative unitarity bounds on the field-space curvature in de Sitter spacetime, using the momentum-space entanglement approach recently proposed by Duaso Pueyo, Goodhew, McCulloch, and Pajer. As an illustration, we perform a perturbative computation of the purity in two-scalar models and compare the resulting unitarity bounds with those obtained via a flat space approximation. In particular, we find that perturbative unitarity imposes an upper bound on the field-space curvature of the Hubble scale order, in addition to a bound analogous to the flat space result. This reflects the thermal nature of de Sitter spacetime. We also discuss generalizations to higher-dimensional field spaces.
@@ 1 Introduction
Notably, EFT becomes more powerful when combined with the concept of ultraviolet (UV) completion and fundamental principles such as unitarity. In particular, the S-matrix unitarity offers a powerful criterion for quantifying the UV cutoff scale and searching for new physics required for UV completion. Historically, the Higgs boson was predicted to restore unitarity in the high-energy scattering of weak bosons Lee et al. (1991, 1977); Dicus and Mathur (1973); Chanowitz and Gaillard (1985). Similarly, string theory, as a UV-complete theory of gravity, emerged from S-matrix theory Veneziano (1968). Moreover, the study of UV completion has in turn revealed that not every EFT is UV completable, leading to nontrivial UV constraints on low-energy effective theories (see, e.g., de Rham et al. (2022) for a review article).
@@
While the S-matrix unitarity offers a fundamental tool to theoretically connect various scales in nature, cosmology is beyond its scope because scattering amplitudes are not well-defined in cosmological backgrounds. To address this limitation, several approaches have been taken so far: A pragmatic approach is to employ the flat space approximation and directly apply implications of the S-matrix unitarity to cosmological models Baumann et al. (2016); Kim et al. (2019); Grall and Melville (2020); Melville and Noller (2020); Kim et al. (2021); Freytsis et al. (2023); Grall and Melville (2022). While this approach has limitations in applicability, the approximation can be justified at least physically as long as our focus is, e.g., on effective couplings generated by the UV dynamics well above the Hubble scale or, in other words, by the dynamics well inside the horizon.
@@
A more challenging direction would be to set up the bootstrap program in cosmology, motivated by recent progress in the S-matrix bootstrap and the conformal bootstrap. Aiming at this ultimate goal, unitarity and (non-)analyticity of cosmological correlators and wavefunctions of the universe have been studied intensively under the slogan of the Cosmological Bootstrap Arkani-Hamed et al. (2020); Baumann et al. (2020, 2021); Arkani-Hamed et al. (2017); Benincasa (2018); Sleight (2020); Sleight and Taronna (2020); Goodhew et al. (2021b); Céspedes et al. (2021); Pajer (2021); Jazayeri et al. (2021); Bonifacio et al. (2021); Melville and Pajer (2021); Goodhew et al. (2021a); Pimentel and Wang (2022); Jazayeri and Renaux-Petel (2022); Qin and Xianyu (2023b); Xianyu and Zhang (2023); Wang et al. (2023); Qin and Xianyu (2023a); Stefanyszyn et al. (2024a); Duaso Pueyo and Pajer (2024); Céspedes et al. (2024); Bzowski et al. (2024); Arkani-Hamed et al. (2023); Grimm et al. (2024); Aoki et al. (2024); Stefanyszyn et al. (2024b); Liu et al. (2025); Goodhew et al. (2024); Ghosh et al. (2025); Lee et al. (2024); Cespedes and Jazayeri (2025); Pimentel and Yang (2025); Stefanyszyn et al. (2025); Qin et al. (2025). There are also alternative attempts to define the notion of the S-matrix itself to cosmological backgrounds Mack (2009); Penedones (2011); Marolf et al. (2013); Melville and Pimentel (2024b); Donath and Pajer (2024); Melville and Pimentel (2024a); Taylor and Zhu (2024); Ferrero and Ripken (2022); Mandal and Banerjee (2021); Spradlin and Volovich (2002); Mei and Mo (2024). These developments motivate further studies toward refinement of the S-matrix unitarity as a guiding principle in cosmology.
@@
Building upon this insight, an interesting approach was proposed recently in Pueyo et al. (2024) to utilize entanglement measures such as purity and entanglement entropy to derive unitarity constraints on cosmological models (see also Colas et al. (2023, 2024); Burgess et al. (2024); Ueda and Tatsumi (2025); Balasubramanian et al. (2012); Aoude et al. (2024); Cheung et al. (2023); Peschanski and Seki (2016); Kowalska and Sessolo (2024); Brahma and Seenivasan (2025); Boutivas et al. (2024) for related developments). A key of this approach is in the fact that interactions induce entanglement between momentum modes, leading to an analogy between entanglement measures and scattering amplitudes. Crucially, this framework is applicable even in curved spacetime as long as the density matrix is well-defined. Based on this approach, perturbative unitarity bounds in inflationary backgrounds were studied in particular.
@@
In this paper, we apply the momentum-space entanglement approach of Pueyo et al. (2024) to derive perturbative unitarity bounds on field-space curvature of nonlinear sigma models, which widely appear for example as effective theories of (pseudo-)Nambu-Goldstone bosons, on a fixed de Sitter background. Unlike the original paper, we study the unitarity bounds on purity without taking the superhorizon limit, which allows us to perform detailed analysis of the bounds that interpolate the flat space analysis and the superhorizon analysis. Interestingly, in addition to a bound similar to the flat space result, we find that the perturbative unitarity gives an upper bound on the field-space curvature of order $H^{-2}$ in terms of the Hubble parameter $H$, which reflects the thermal nature of de Sitter spacetime.
@@
This paper is organized as follows: In Sec. 2, we review the momentum-space entanglement approach to perturbative unitarity proposed in Pueyo et al. (2024). In particular, we introduce a perturbative formula of purity and its unitarity condition. In Sec. 3, we study the perturbative unitarity bounds in flat spacetime and show that the UV cutoff is set by the field-space curvature, similarly to the bounds obtained from scattering amplitudes. In Sec. 4, we extend the analysis to de Sitter spacetime. In addition to a bound similar to the flat space result, we find an upper bound on the field-space curvature at the order of the Hubble scale. We conclude our analysis in Sec. 5. Technical details are collected in the Appendices.
@@
Throughout the paper, we adopt the metric signature (,+,+,+)(-,+,+,+), Greek letters $\mu,\nu,\cdots$ to denote spacetime indices, and shorthand notations for the integration measure,
@@ 2 Perturbative unitarity bounds from purity: a brief review
This section gives a brief review of the momentum-space entanglement approach to perturbative unitarity bounds proposed in Pueyo et al. (2024). In particular, we consider purity that quantifies the entanglement between a system of interest and its complement (environment). In Sec. 2.1, we first introduce the concept of purity in quantum field theory (QFT) and briefly explain how it can be evaluated using the wavefunction representation. Then, in Sec. 2.2, we consider multi-scalar models with nonzero field-space curvature and provide a concrete formula for the purity, which is used in the following sections.
@@ 2.1 Momentum-space entanglement and purity in EFT
Consider a quantum system in a pure state represented by a density matrix [^1] In Pueyo et al. (2024), the density matrix was defined without being canonically normalized to carefully discuss regularization for the continuum and infinite-volume limit. The density matrix here should also be understood under the same regularization, even though we assume canonical normalization for visual clarity.,
[…]
@@
$\rho$ $|\displaystyle=\ket{\Omega}\bra{\Omega}$
$|\displaystyle=\int\mathcal{D}\phi\mathcal{D}\phi^{\prime}\ket{\phi}\braket{\phi|\Omega}\braket{\Omega|\phi^{\prime}}\bra{\phi^{\prime}}$
$|,\displaystyle=\int\mathcal{D}\phi\mathcal{D}\phi^{\prime}\rho_{\phi\phi^{\prime}}\ket{\phi}\bra{\phi^{\prime}}\,$, (5)
where we define the components of the density matrix as $\rho_{\phi\phi^{\prime}}\coloneqq\braket{\phi|\Omega}\braket{\Omega|\phi^{\prime}}$. Here and in the rest of this subsection, we focus on a single real scalar model for illustration, but its extension to general setups is straightforward.
@@
Its violation signals breakdown of the EFT, hence we can use the purity bound (9) to identify the maximum energy range of validity of the EFT [^2] Alternatively, if the energy scale of interest is specified and the cutoff scales $\Lambda_{\text{IR}}$ and $\Lambda_{\text{UV}}$ are given, one can interpret the bounds (9) as consistency conditions on the model parameters such as the particle spectrum and the coupling constants. .
@@
[…]
$\Psi[\phi]\coloneqq\braket{\phi|\Omega}\,.$ (10)
@@
[…]
@@ 2.2 Perturbative formula for purity
$S$ $=\int{\mathrm{d}}^{4}x\sqrt{-g}\left[-\frac{1}{2}\sum_{I}\Big((\partial_{\mu}\phi^{I})^{2}+m_{I}^{2}(\phi^{I})^{2}\Big)+\frac{1}{2}\sum_{I,J,K,L}C_{IJKL}\phi^{K}\phi^{L}\partial_{\mu}\phi^{I}\partial^{\mu}\phi^{J}+\cdots\right]\,$, (15)
where the dots stand for higher order terms in $\phi^{I}$ and they are irrelevant in the following analysis. We study this model in homogeneous and isotropic spacetime,
@@
[…]
@@
$\hskip 60.0pt\times\left(\partial_{\eta}K_{\phi^{I}}(k_{1};\eta)\partial_{\eta}K_{\phi^{J}}(k_{2};\eta)+({\bm{k}}_{1}\cdot{\bm{k}}_{2})K_{\phi^{I}}(k_{1};\eta)K_{\phi^{J}}(k_{2};\eta)\right)\Big]$
+ $\quad+\text{ ($23$ perm.)}\,$, (19)
@@
Finally, we provide a formula for the purity. Let us choose the Fourier modes $\phi^{\bar{I}}_{{\bm{p}}}$ and $\phi^{\bar{I}}_{-{\bm{p}}}$ of the species label $\bar{I}$ as the system of interest. Then, from the definition (6)-(7) and the wavefunction representation of the density matrix (11), the purity can be evaluated at the tree level as [^3] See the original paper Pueyo et al. (2024) for details of the diagrammatic method for the purity computation.
@@
[…]
@@ 3 Perturbative unitarity bounds in flat spacetime
We begin with the flat spacetime, for which perturbative unitarity bounds on the field-space curvature were already studied using scattering amplitudes (see, e.g., Nagai et al. (2019)). To be concrete, this section focuses on the following two-scalar model as a simple case of (15):
@@ 3.2.1 Bounds on $\gamma_{\phi}$
First, we consider $\gamma_{\phi}$, choosing the Fourier modes $\phi_{\bm{p}}$ and $\phi_{-{\bm{p}}}$ as the system. From the formula (21), the corresponding purity reads [^5] Thanks to the rotational symmetry, the momentum-dependence of the purity in our setup is only through the amplitude $p=|{\bm{p}}|$ of the momentum $\bm{p}$} of the system modes. Also, it is convenient for visual clarity to suppress its cutoff-dependence. Hence, we employ the notation $\gamma_{\phi^{\bar{I}}}(p)$ and $I_{\phi^{\bar{I}}}(p)$ for $\gamma_{\phi^{\bar{I}}}(\Lambda_{\text{IR}},\Lambda_{\text{UV}},{\bm{p}})$ and $I_{\phi^{\bar{I}}}(\Lambda_{\text{IR}},\Lambda_{\text{UV}},{\bm{p}})$ in the rest of the paper.
@@ 4.1 Illustrative example: $m_{\phi}=0$ and $m_{\sigma}=\sqrt{2}H$
$\quad\times\bigg[k_{1}k_{2}\Big(1+(1-ik_{T}\eta_{*})^{2}\Big)(1-\cos\theta)-\Big((k_{1}+k_{2})\left(1-ik_{T}\eta_{*}\right)+k_{T}\Big)k_{T}\cos\theta\bigg]\,$, (50)
@@
Note that the UV cutoff scale has to be well above the Hubble scale $\Lambda_{\text{UV}}\gtrsim H$, otherwise the horizon-scale dynamics cannot be captured. Below, we present the bounds on $\gamma_{\phi}$ and $\gamma_{\sigma}$, and their implications in order.